OpenStudy (anonymous):

Factor completely: a) x^3-4x^2-11x+30 b) 2x^3+3x^2-1

OpenStudy (anonymous):

use factor and remainder theorems..

OpenStudy (anonymous):

x^3-4x^2-11x+30 =x^3-2x^22x^2+4x-15x+30 =x^2(x-2) -2x(x-2) -15(x-2) =(x-2)(x^2-2x-15) =(x-2)(x^2-5x+3x-15) =(x-2)( x(x-5) +3(x-5)) =(x-2)(x+3)(x-5)

OpenStudy (anonymous):

2x^3+3x^2-1 =2x^3+3x^2-3+2 =2(x^3+1) +3(x^2-1) =2(x+1)(x^2-x+1)+3(x+1)(x-1) =(x+1)(2x^2 -2x+2+3x-3) =(x+1)(2x^2 -x -5)